. The selling price for each item of a product is $50 and Total cost is given by


๐ถ(๐‘ฅ) = 10๐‘ฅ + 18,000

Where ๐‘ฅ is the number of items.

a) Write revenue function.

b) Find ๐ถ(100) and write a sentence that explains itโ€™s meaning.

c) Find number of units that gives break-even point.


1
Expert's answer
2021-04-15T06:49:39-0400

a)

"R(x)=p\cdot x=50x"

b)


"C(100)=10(100)+18000=19000"

The total cost of producing 100 units is $19,000.


c)

"P(x)=0=>50x-(10x+18000)=0"

"40x=18000"

"x=450"